# Cointegration test in time series analysis by global optimisation

### v2

@@ Abstract.
Time Series Analysis has been given a great amount of study in which many useful tests were developed. The phenomenal work of Engle and Granger [1] in 1987 and Johansen [2] in 1988 has paved the way for the most commonly used cointegration tests so far. Even though cointegrating relationships focus on long-term behaviour and correlation of multiple nonstationary time series, oftentimes we encounter statistical data with limited sample sizes and other information. Thus other tests with empirical advantages may also be of considerable importance.
In this paper, we provide an optimisation approach motivated by the Blind Source Separation, or also known as Independent Component Analysis, for cointegration between financial time series. Two methods for cointegration tests are introduced, namely decorrelation for the bivariate case and maximisation of nongaussianity for higher-dimensions. We highlight the empirical preponderances of independent components and also the computational simplicity, compared to common practices of cointegration such as the Johansen’s Cointegration Test [3]. The advantages of our methods, especially the better performances in limited sample size, enable a wider range of application and accessibility for researchers and practitioners to identify cointegrating relationships.
Key words and phrases:
Cointegration; mean-reverting; convergence trading; financial time series; blind source separation.
@@ 1. Introduction
The structure of this paper is as follows, section 2 introduces preliminary knowledge about multivariate time series and the basic algorithms behind blind source separation, specifically the methodology of decorrelation and maximization of nongaussianity. [11]. Furthermore, the Johansen’s cointegration test would also be introduced for further reference and comparison with our own algorithm (see Appendix). In section 3, we discuss the theoretical background and convergence properties of the methods. Moving on to section 4, the paper provides practical applications of the methods using both simulated series and oil prices as a real-life example. Section 4 also brings out the empirical advantages and drawbacks with comparisons within our methods and Johansen’s method. In Section 5, a comprehensive conclusion is provided.
@@ 2.1. Vector Autoregressive Models
By introducing the Back-Shift Operator (i.e. a matrix B such that $\mathbf{s}_{t}=B\mathbf{s}_{t-1}$ for any $t$), we see that a VAR(p) process can be transformed into
@@
Now consider $\mathbf{x}_{t}:=(\mathbf{\tilde{s}}_{t-p+1},\mathbf{\tilde{s}}_{t-p+2},\ldots,\mathbf{\tilde{s}}_{t})\in\mathbb{R}^{k}\times\mathbb{R}^{p}$ and $\bm{\delta}_{t}:=(\mathbf{0},\mathbf{0},\ldots,\bm{\varepsilon}_{t})\in\mathbb{R}^{k}\times\mathbb{R}^{p}$. A VAR(p) model is equivalent to
@@
$\Phi_{*}=\begin{bmatrix}\mathbf{0}&I&\mathbf{0}&\mathbf{0}&\ldots&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&I&\mathbf{0}&\ldots&\mathbf{0}\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\ldots&I\\ \Phi_{p}&\Phi_{p-1}&\Phi_{p-2}&\Phi_{p-3}&\ldots&\Phi_{1}\end{bmatrix}.$
@@
$\mathbf{x}_{t}-\bm{\mu}=\bm{\delta}_{t}+\sum_{i=1}^{p}{\Phi_{*}^{i}}\bm{\delta}_{t-i}.$
@@ 2.2. The blind source separation
where we assume that $\mathbf{x},\mathbf{s}\in\mathbb{R}^{n}$ and $A\in\mathbb{R}^{n\times n}$. In a practical problem setting, $\mathbf{x}$ is known in terms of observed data, we then separate the vector $\mathbf{s}$ from the mixing matrix $A$, which is assumed to be nonsingular.
@@ 2.2.1. The decorrelation method
(2.2) $\begin{cases}v_{1t}=(a_{11}a_{22}-a_{12}a_{21})s_{1t}=det(A)s_{1t},\\ v_{2t}=\det(A)s_{2t}.\end{cases}$
@@ 2.2.2. Maximisation of nongaussianity
$kurt(y)=\operatorname{\mathbb{E}}[y^{4}]-3(\operatorname{\mathbb{E}}[y^{2}])^{2}.$
@@
2:$\mathbf{w}_{k+1}\leftarrow\operatorname{\mathbb{E}}[\mathbf{x}^{T}f(\mathbf{w}_{k}^{T}\mathbf{x})]-\operatorname{\mathbb{E}}[f^{\prime}(\mathbf{w}_{k}^{T}\mathbf{x})]\mathbf{w}_{k}$
3:$\mathbf{w}_{k}\leftarrow\frac{\mathbf{w}_{k}}{\lVert\mathbf{w}_{k}\rVert}$triangleright If not converge, go back to 2
@@ 3.1. Problem setting
Let $s_{1t}$ and $s_{2t}$ be two non-stationary time series. Let $r_{t}$ be an ARIMA$(p,q,d)$ process. Let $\varepsilon_{t}$ be a white noise series. Consider the following system
@@
Upon solving the quadratic equations we get the mixing matrix $B$ and hence the recovered signal series $r_{t}$ and $\epsilon_{t}$.
@@
Since $S(t)$ is nonstationary, if we have $\beta_{1}a_{11}+\beta_{2}a_{21}=0$, then
@@
$A\bm{\beta}=\begin{bmatrix}a_{11}&a_{21}\\ a_{12}&a_{22}\end{bmatrix}\begin{bmatrix}\beta_{1}\\ \beta_{2}\end{bmatrix}=\begin{bmatrix}0\\ *\end{bmatrix}.$
@@
(3.3) $\operatorname*{arg\,min}_{\beta_{1},\beta_{2}}$ $\operatorname{\mathbb{E}}[(\beta_{1}s_{1t}+\beta_{2}s_{2t})^{2}]$
@@
For a generalised version where one has multiple non-stationary processes, if we have $(n-1)$ mutually uncorrelated nonstationary autoregressive integrated moving average (ARIMA) processes $r_{1t},r_{2t},\ldots,r_{n-1,t}$ with an order of integration $I(d)$, $d\geq 1$ [13, 12, 14]. In statistics, the order of integration, denoted $I(d)$, means the minimum number of differences required to obtain a covariance-stationary series. Let $\varepsilon_{t}$ be an independent Gaussian process (stationary series). Suppose there is a nonsingular matrix $A=(a_{ij})_{n\times n}$ such that $\mathbf{s}_{t}=A\begin{pmatrix}\mathbf{r}_{t}\\ \varepsilon_{t}\end{pmatrix}$, where $\mathbf{s}_{t}=(s_{1t},\ldots,s_{nt})^{T}\in\mathbb{R}^{n}$ and $\mathbf{r}_{t}=(r_{1t},\ldots,r_{n-1,t})^{T}\in\mathbb{R}^{n-1}$.
@@
(3.5) $=$ $\operatorname*{arg\,min}_{\bm{\beta}}\sum_{j=1}^{n-1}{\left(\sum_{i=1}^{n}{\beta_{i}a_{ij}}\right)^{2}\operatorname{\mathbb{E}}[r_{jt}^{2}]}+\sum_{i=1}^{n}{\beta_{i}^{2}a_{in}^{2}\operatorname{\mathbb{E}}[\varepsilon_{t}^{2}]}$
@@
3:$\mathbf{w}_{k+1}^{(i)}\leftarrow\operatorname{\mathbb{E}}[\mathbf{x}^{T}f({\mathbf{w}_{k}^{(i)}}^{T}\mathbf{x})]-\operatorname{\mathbb{E}}[f^{\prime}({\mathbf{w}_{k}^{(i)}}^{T}\mathbf{x})]\mathbf{w}_{k}^{(i)}$
4:$\mathbf{w}^{(i)}\leftarrow\mathbf{w}^{(i)}-\sum_{j=1}^{i-1}\left({\mathbf{w}^{(j)}}^{T}\mathbf{w}^{(i)}\right)\mathbf{w}^{(j)}$
5:$\mathbf{w}_{k}^{(i)}\leftarrow\frac{\mathbf{w}_{k}^{(i)}}{\lVert\mathbf{w}_{k}^{(i)}\rVert}$triangleright If not converge, go back to 3
6:$p\leftarrow p+1$triangleright If $p<n$, go back to 2
@@
$\mathbf{w}^{(i)}\leftarrow\mathbf{w}^{(i)}-\sum_{j=1}^{i-1}\left({\mathbf{w}^{(j)}}^{T}\mathbf{w}^{(i)}\right)\mathbf{w}^{(j)}.$
@@ 3.4. Convergence and stability analysis
It is important to highlight that in practical applications, computing kurtosis or negentropy as a metric of nongaussianity can be challenging, necessitating the use of reliable and precise approximations to streamline the calculation process. Through the selection of suitable functions $F_{i}:\mathbb{R}\to\mathbb{R}$ for $i=1$, $2$, it is possible to derive an effective approximation of negentropy, as demonstrated below:
@@
where $v$ is a standardised (normal) random variable. The constants $k_{i}$ for $i=1$, $2$ are given by $k_{i}^{2}=\frac{1}{2\delta_{i}^{2}}$ where
[…]
@@
$\mathbf{w}\leftarrow\operatorname{\mathbb{E}}[\mathbf{s}f(\mathbf{w}^{T}\mathbf{s})].$
@@
(3.10) $\mathbf{w}$ $=\operatorname{\mathbb{E}}[\mathbf{s}f(\mathbf{w}^{T}\mathbf{s})],$
(3.11) $(1+\alpha)\mathbf{w}$ $=\operatorname{\mathbb{E}}[\mathbf{s}f(\mathbf{w}^{T}\mathbf{s})]+\alpha\mathbf{w}.$
@@
(3.12) $\operatorname{\mathbb{E}}[\mathbf{s}f\left(\mathbf{w}^{T}\mathbf{s}\right)]+\lambda\mathbf{w}=\mathbf{0}.$
@@
$\left\lVert\mathbf{x}_{k}-\mathbf{x}^{*}\right\rVert_{\infty}\leq\frac{K^{k}}{1-K}\left\lVert\mathbf{x}_{1}-\mathbf{x}_{0}\right\lVert_{\infty}.$
@@
    $\frac{\left\lvert\sum_{t=1}^{T}r_{it}\right\lvert}{T}\leq c_{0}.$
@@
    $\frac{\left\lvert\sum_{t=1}^{T}g(r_{it})f(r_{it})-\sum_{t=1}^{T}g(r_{it})\sum_{t=1}^{T}f(r_{it})\right\lvert}{T}\leq c_{T}.$
@@
$\operatorname{\mathbb{E}}\left[\sum_{j\neq i}f^{\prime}(\eta r_{it})r_{jt}^{2}-f(\eta r_{it})r_{it}\right]\begin{cases}>0&\text{for maxima,}\\ <0&\text{for minima.}\end{cases}$
@@
$+\frac{1}{2}\sum_{j\neq i}\sum_{l\neq i,j}\operatorname{\mathbb{E}}[f^{\prime}(\eta r_{it})r_{jt}r_{lt}]\epsilon_{j}\epsilon_{l}+\frac{1}{2}\sum_{j\neq i}\operatorname{\mathbb{E}}\left[f^{\prime}(\eta r_{it})r_{jt}^{2}\right]\epsilon_{j}^{2}+o(\lVert\bm{\epsilon}\rVert^{2}).$
@@
$+\frac{1}{2}\sum_{j\neq i}\sum_{l\neq i,j}\operatorname{\mathbb{E}}[f^{\prime}(\eta r_{it})]c_{T}^{2}\epsilon_{j}\epsilon_{l}+\frac{1}{2}\sum_{j\neq i}\operatorname{\mathbb{E}}\left[f^{\prime}(\eta r_{it})r_{jt}^{2}\right]\epsilon_{j}^{2}+o(\lVert\bm{\epsilon}\rVert^{2}).$
@@
$\leq H(\eta\mathbf{e}_{i})+\frac{1}{2}\left(\sum_{j\neq i}\operatorname{\mathbb{E}}\left[f^{\prime}(\eta r_{it})r_{jt}^{2}\right]-\operatorname{\mathbb{E}}\left[f(\eta r_{it})r_{it}\right]\right)\sum_{j\neq i}\epsilon_{j}^{2}+o(\lVert\bm{\epsilon}\rVert^{2}),$

### v3

# Estimation of Cointegration Vectors in Time Series via Global Optimisation

@@ Abstract.
Time Series Analysis has been given a great amount of study in which many useful tests were developed. The phenomenal work of Engle and Granger [1] in 1987 and Johansen [2] in 1988 has paved the way for the most commonly used cointegration tests so far. Even though cointegrating relationships focus on long-term behaviour and correlation of multiple nonstationary time series, oftentimes we encounter statistical data with limited sample sizes and other information. Thus other tests with empirical advantages may also be of considerable importance. In this paper, we provide an optimisation approach motivated by the Blind Source Separation, or also known as Independent Component Analysis, for cointegration between financial time series. Two methods for cointegration tests are introduced, namely decorrelation for the bivariate case and maximisation of nongaussianity for higher-dimensions. We highlight the empirical preponderances of independent components and also the computational simplicity, compared to common practices of cointegration such as the Johansen’s Cointegration Test [3]. The advantages of our methods, especially the better performances in limited sample size, enable a wider range of application and accessibility for researchers and practitioners to identify cointegrating relationships.
@@ 1. Introduction
The structure of this paper is as follows, section 2 introduces preliminary knowledge about multivariate time series and the basic algorithms behind blind source separation, specifically the methodology of decorrelation and maximization of nongaussianity. [11]. Furthermore, the Johansen’s cointegration test would also be introduced for further reference and comparison with our own algorithm (see Appendix). In section 3, we discuss the theoretical background and convergence properties of the methods. Moving on to section 4, the paper provides practical applications of the methods using both simulated series and oil prices as a real-life example. Section 4 also brings out the empirical advantages and drawbacks with comparisons within our methods and Johansen’s method.
@@ 2.1. Vector Autoregressive Models
By introducing the Back-Shift Operator, we see that a VAR(p) process can be transformed into
@@
Now consider $\mathbf{x}_{t}:=(\mathbf{s}_{t-p+1},\mathbf{s}_{t-p+2},\ldots,\mathbf{s}_{t})\in\mathbb{R}^{k}\times\mathbb{R}^{p}$ and $\bm{\delta}_{t}:=(\mathbf{0},\mathbf{0},\ldots,\bm{\varepsilon}_{t})\in\mathbb{R}^{k}\times\mathbb{R}^{p}$. A VAR(p) model is equivalent to
@@
$\Phi_{*}=\begin{bmatrix}\mathbf{0}&I&\mathbf{0}&\mathbf{0}&\ldots&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&I&\mathbf{0}&\ldots&\mathbf{0}\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\ldots&I\\ \Phi_{p}&\Phi_{p-1}&\Phi_{p-2}&\Phi_{p-3}&\ldots&\Phi_{1}\end{bmatrix}$
@@
$\mathbf{x}_{t}-\bm{\mu}=\bm{\delta}_{t}+\sum_{i=1}^{p}{\Phi_{*}^{i}}\bm{\delta}_{t-i}$
@@ 2.2. The blind source separation
where we assume that $\mathbf{x},\mathbf{s}\in\mathbb{R}^{n}$ and $A\in\mathbb{R}^{n\times n}$. In a practical problem setting, $\mathbf{x}$ is known in terms of observed data, we then separate the vector $\mathbf{s}$, which is assumed to be consisting of independent components, from the mixing matrix $A$, which is assumed to be nonsingular.
@@ 2.2.1. The decorrelation method
(2.2) $\begin{cases}v_{1t}=(a_{11}a_{22}-a_{12}a_{21})s_{1t}=det(A)s_{1t},\\ v_{2t}=\det(A)s_{2t}.\end{cases}\$
@@ 2.2.2. Maximisation of nongaussianity
$kurt(y)=\operatorname{\mathbb{E}}[y^{4}]-3(\operatorname{\mathbb{E}}[y^{2}])^{2}$
@@
[…]
@@ 3.1. Problem setting
Let $s_{1t}$ and $s_{2t}$ be two non-stationary time series. Let $r_{t}$ be an ARIMA(p,d,q)(p,d,q) process. Let $\varepsilon_{t}$ be a white noise series. Consider the following system
@@
Upon solving the quadratic equations we get the mixing matrix $A$ and hence the recovered signal series $r_{t}$ and $\epsilon_{t}$.
@@
Since $r_{t}$ is nonstationary, if we have $\beta_{1}a_{11}+\beta_{2}a_{21}=0$, then
@@
$A\bm{\beta}=\begin{bmatrix}a_{11}&a_{21}\\ a_{12}&a_{22}\end{bmatrix}\begin{bmatrix}\beta_{1}\\ \beta_{2}\end{bmatrix}=\begin{bmatrix}0\\ *\end{bmatrix}$
@@
(3.3) $\operatorname*{arg\,min}_{\beta_{1},\beta_{2}}$ $\operatorname{\mathbb{E}}[(\beta_{1}s_{1t}+\beta_{2}s_{2t})^{2}].$
@@
For a generalised version where one has multiple non-stationary processes, if we have (n1)(n-1) mutually uncorrelated nonstationary autoregressive integrated moving average (ARIMA) processes $r_{1t},r_{2t},\ldots,r_{n-1,t}$ with an order of integration $I(d)$, $d\geq 1$ [13, 12, 14]. In statistics, the order of integration, denoted $I(d)$, means the minimum number of differences required to obtain a covariance-stationary series. Let $\varepsilon_{t}$ be an independent Gaussian process (stationary series). Suppose there is a nonsingular matrix $A=(a_{ij})_{n\times n}$ such that $\mathbf{s}_{t}=A\begin{pmatrix}\mathbf{r}_{t}\\ \varepsilon_{t}\end{pmatrix}$, where $\mathbf{s}_{t}=(s_{1t},\ldots,s_{nt})^{T}\in\mathbb{R}^{n}$ and $\mathbf{r}_{t}=(r_{1t},\ldots,r_{n-1,t})^{T}\in\mathbb{R}^{n-1}$.
@@
(3.5) $=$ $\operatorname*{arg\,min}_{\bm{\beta}}\sum_{j=1}^{n-1}{\left(\sum_{i=1}^{n}{\beta_{i}a_{ij}}\right)^{2}\operatorname{\mathbb{E}}[r_{jt}^{2}]}+\sum_{i=1}^{n}{\beta_{i}^{2}a_{in}^{2}\operatorname{\mathbb{E}}[\varepsilon_{t}^{2}]}.$
@@
[…]
@@
$\mathbf{w}^{(i)}\leftarrow\mathbf{w}^{(i)}-\sum_{j=1}^{i-1}\left({\mathbf{w}^{(j)}}^{T}\mathbf{w}^{(i)}\right)\mathbf{w}^{(j)}$
@@ 3.4. Convergence and stability analysis
It is important to highlight that in practical applications, computing kurtosis or negentropy as a metric of nongaussianity can be challenging, necessitating the use of reliable and precise approximations to streamline the calculation process. Through the selection of suitable functions $F_{i}:\mathbb{R}\to\mathbb{R}$ for $i=1$, 22, it is possible to derive an effective approximation of negentropy, as demonstrated below:
@@
where $v$ is a standardised (normal) random variable. The constants $k_{i}$ for $i=1$, 22 are given by $k_{i}^{2}=\frac{1}{2\delta_{i}^{2}}$ where
$\delta_{1}^{2}$ $=\int\phi(\xi)\left(F_{1}(\xi)\right)^{2}\ d\xi-\left(\int\phi(\xi)F_{1}(\xi)\xi\ d\xi\right)^{2}$
$\delta_{2}^{2}$ $=\int\phi(\xi)(F_{2}(\xi))^{2}\ d\xi-\left(\int\phi(\xi)F_{2}(\xi)\ d\xi\right)^{2}-\frac{1}{2}\left(\int\phi(\xi)F_{2}(\xi)\ d\xi-\int\phi(\xi)F_{2}(\xi)\xi^{2}\ d\xi\right)^{2}$
@@
$\mathbf{w}\leftarrow\operatorname{\mathbb{E}}[\mathbf{s}f(\mathbf{w}^{T}\mathbf{s})]$
@@
(3.10) $\mathbf{w}$ $=\operatorname{\mathbb{E}}[\mathbf{s}f(\mathbf{w}^{T}\mathbf{s})]$
(3.11) $(1+\alpha)\mathbf{w}$ $=\operatorname{\mathbb{E}}[\mathbf{s}f(\mathbf{w}^{T}\mathbf{s})]+\alpha\mathbf{w}$
@@
(3.12) $\operatorname{\mathbb{E}}[\mathbf{s}f\left(\mathbf{w}^{T}\mathbf{s}\right)]+\lambda\mathbf{w}=\mathbf{0}$
@@
$\left\lVert\mathbf{x}_{k}-\mathbf{x}^{*}\right\rVert_{\infty}\leq\frac{K^{k}}{1-K}\left\lVert\mathbf{x}_{1}-\mathbf{x}_{0}\right\lVert_{\infty}$
@@
    $\frac{\left\lvert\sum_{t=1}^{T}r_{it}\right\lvert}{T}\leq c_{0}$
@@
    $\frac{\left\lvert\sum_{t=1}^{T}g(r_{it})f(r_{it})-\sum_{t=1}^{T}g(r_{it})\sum_{t=1}^{T}f(r_{it})\right\lvert}{T}\leq c_{T}$
@@
$\operatorname{\mathbb{E}}\left[\sum_{j\neq i}f^{\prime}(\eta r_{it})r_{jt}^{2}-f(\eta r_{it})r_{it}\right]\begin{cases}>0&\text{for maxima}\\ <0&\text{for minima}\end{cases}$
@@
$+\frac{1}{2}\sum_{j\neq i}\sum_{l\neq i,j}\operatorname{\mathbb{E}}[f^{\prime}(\eta r_{it})r_{jt}r_{lt}]\epsilon_{j}\epsilon_{l}+\frac{1}{2}\sum_{j\neq i}\operatorname{\mathbb{E}}\left[f^{\prime}(\eta r_{it})r_{jt}^{2}\right]\epsilon_{j}^{2}+o(\lVert\bm{\epsilon}\rVert^{2})$
@@
$+\frac{1}{2}\sum_{j\neq i}\sum_{l\neq i,j}\operatorname{\mathbb{E}}[f^{\prime}(\eta r_{it})]c_{T}^{2}\epsilon_{j}\epsilon_{l}+\frac{1}{2}\sum_{j\neq i}\operatorname{\mathbb{E}}\left[f^{\prime}(\eta r_{it})r_{jt}^{2}\right]\epsilon_{j}^{2}+o(\lVert\bm{\epsilon}\rVert^{2})$
@@
$\leq H(\eta\mathbf{e}_{i})+\frac{1}{2}\left(\sum_{j\neq i}\operatorname{\mathbb{E}}\left[f^{\prime}(\eta r_{it})r_{jt}^{2}\right]-\operatorname{\mathbb{E}}\left[f(\eta r_{it})r_{it}\right]\right)\sum_{j\neq i}\epsilon_{j}^{2}+o(\lVert\bm{\epsilon}\rVert^{2})$
@@
The maximization of nongaussianity approach naturally handles higher-dimensional cointegration systems by separating all independent components simultaneously. Each recovered component is tested for stationarity using the Augmented Dickey-Fuller test, with the stationary components corresponding to cointegrating relationships. The cointegration vectors are directly obtained from the rows of the demixing matrix associated with stationary components.
However, notice that the application of the ADF test to components recovered via an optimization procedure may lead to non-standard asymptotic distributions of the test statistic, a problem analogous to that encountered when testing OLS regression residuals for cointegration [1]. The critical values used in this paper should therefore be interpreted with caution. A more rigorous analysis would require the simulation of specific critical values for this BSS-based procedure, which we leave for future research.
Remark.
The optimization in (3.7) relies critically on the non-Gaussianity of the stationary components. If all stationary components follow Gaussian distributions, the negentropy measure $J(\mathbf{y})$ becomes zero, and the objective function $\mathscr{G}(\cdot)$ fails to identify meaningful independent components. This limitation arises because Gaussian distributions are rotationally invariant under orthogonal transformations, making separation impossible.
However, this constraint does not invalidate our method in practical applications. Firstly, financial/economic time series frequently exhibit non-Gaussian features (e.g., heavy tails from volatility clustering). Secondly, cointegration residuals often demonstrate non-normality due to structural breaks or heteroskedasticity. Lastly, the decorrelation method (Section 3.2) remains valid for Gaussian systems when second-order statistics suffice. In scenarios with suspected Gaussianity, we recommend using Johansen’s test or verifying non-normality via Jarque-Bera tests prior to applying our method.